ne^2 tau /m = 1/RA derivationHow this formula came?

Grade:10

1 Answers

Mohammed Kasib Siddiqui
5 years ago
We know F= qE
for electrons-
F=eE
ma=eE
a=eE/m
v/t=eE/m            (v=at)  ( t= tau or relaxation time)
v= eE t/m
i/(nea)= eV t/ml     (v= drift velocity) (E=V/l)
i/V=R= ml/(ne^2 t A)
If length of conductor is 1m then l=1
Hence..
ne^2t/m= 1/RA ans.
 
Sorry for the late answer